# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin-Planck principle

### v5

@@ Abstract
We derive the law of entropy non-decrease directly from the Kelvin-Planck principle for simple and compound systems without using the Clausius inequality. A key of the derivation is a formulation of entropy in terms of work by a Carnot machine operating between a system and a single heat reservoir at fixed temperature, which is equivalent to Clausius entropy based on heat and Gyftopoulos-Beretta entropy based on work. We also show that we can characterize entropy as the thermodynamic cost to create a system’s state from a reference state.
I Introduction
How information of irreversibility is encoded into entropy as a state quantity? This would be one of the most difficult points for those who learn thermodynamics first to grasp. This is also true for physicists today, and the famous paper by Lieb and Yngvason was aimed to elucidate the nature of entropy in terms of adiabatic accessibility in their axiomatic approach to thermodynamics [1, 2, 3], where neither statistical mechanics nor heat engines are used [3] (see also Ref. [4]).
Here, let us consider a conventional approach to thermodynamics. Consider a thermally isolated system in an equilibrium state $A_{1}$. After an arbitrary adiabatic process that changes the system’s state to another equilibrium state $A_{2}$, the system’s entropy $S$ never decreases before and after the change:
@@ I Introduction
where $S_{0}$ is the entropy at a reference state $A_{0}$, $\delta Q$ denotes heat supplied to the system, and $T_{R}$ denotes the temperature of a heat reservoir. During the process from $A_{0}$ to $A_{1}$, the temperature $T_{R}$ of the heat reservoir that contacts the system may change [5]. Equation (2) is the formula derived by Clausius. Usually, in standard textbooks on thermodynamics such as Ref. [6], after introducing Carnot cycle and Carnot efficiency as the maximum efficiency of heat engines, Eq. (1) is derived by using the Clausius inequality for a cyclic process:
@@
where $W_{\rm cyc}$ denotes the work extracted from a cycle starting from an equilibrium state, and the equality applies to a reversible cycle. Deriving the Clausius inequality from the Kelvin-Planck principle relies on a complicated thermodynamic process that tactically involves multiple (or infinitely many) heat reservoirs at different temperatures and multiple (or infinitely many) Carnot heat engines [6]. This is because the Kelvin-Planck principle is originally applied to a process with a single heat reservoir at fixed temperature, and a complicated setup is necessary for it to be applied to a case with multiple (or infinitely many) heat reservoirs. Moreover, the law of entropy non-decrease Eq. (1) is derived as a special case of the Clausius inequality Eq. (3) (i.e., the case of $\delta Q=0$ in Eq. (4)) and it is thus derived indirectly from the Kelvin-Planck principle.
@@
In this paper, we show that we can derive the law of entropy non-decrease Eq. (1) directly from the Kelvin-Planck principle Eq. (5) without using the Clausius inequality Eq. (3). To this end, we give a new formulation of entropy of a substance in terms of the work of a Carnot machine operating between the substance and a single heat reservoir at fixed temperature. Because we will assume only a single heat reservoir at fixed temperature in our derivation, it apparently makes it possible to apply the Kelvin-Planck principle to show the law of entropy non-decrease. We discuss the relationship between our entropy and the Clausius entropy (2) based on heat or the Gyftopoulos-Beretta entropy [7] based on work. Moreover, we also show that entropy may be characterized by the thermodynamic cost to create a system’s state from a reference state.
II Main results
@@
We derive the law of entropy non-decrease Eq. (1) from the Kelvin-Planck principle Eq. (5). After the formulation of our entropy in terms of work by a Carnot machine in Sec. II.1, we give the derivation for simple systems in Sec II.2 and compound systems in Sec. II.3. Here, a simple system is a system in an equilibrium state that can be specified with the system’s temperature $T$ and volume $V$. A compound system is a system composed of simple systems.
II.1 Formulation of entropy as state quantity
@@
Figure 1: A reversible path $\gamma$ from $A_{0}$ to $A_{1}$ and a reversible path $\gamma^{\prime}$ from $A_{1}$ to $A_{0}$ on the temperature-volume ($T$$V$) plane. They consist of the reversible adiabatic processes (arrows along the adiabatic curves) and the reversible temperature-changing processes with a Carnot machine operating between a system at constant volume and a heat reservoir at temperature $T_{0}$ (vertical arrows).
@@
We introduce our entropy as a state quantity in the following way. Consider a simple system initially in an equilibrium state $A_{0}$ specified with $(T_{0},V_{0})$ with $T_{0}$ and $V_{0}$ the system’s initial temperature and volume, respectively. Consider a reversible path $\gamma$ from the equilibrium state $A_{0}$ to another equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$ and a reversible path $\gamma^{\prime}$ from the state $A_{1}$ to $A_{0}$ (Fig. 1). Any reversible path $\gamma$ from the equilibrium state $A_{0}$ to another one $A_{1}$ can be realized by a combination of (i) a reversible adiabatic process (the arrow along the adiabatic curve in Fig. 1) and (ii) a reversible temperature-changing process with a Carnot machine operating between the system at constant volume and a single heat reservoir at fixed temperature $T_{o}$ to raise or lower the system’s temperature (the vertical arrow in Fig. 1) by regarding the system as a finite-sized heat reservoir [5]; any reversible path $\gamma$ may be decomposed into $N$-reversible processes $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$ with $\gamma_{i}$ being the $i$-th reversible process, which is either of the above two thermodynamic processes (i) and (ii). In particular, when the path is continuous, it may be realized by a combination of the processes (i) and (ii) by taking the limit of $N\to\infty$. It should be noted that there is no need for the system’s temperature to match the fixed temperature $T_{o}$ of the heat reservoir during the process including the initial and final states $A_{0}$ and $A_{1}$.
@@
where $-\gamma^{\prime}$ denotes the reversed path of $\gamma^{\prime}$. Because $\gamma$ and $-\gamma^{\prime}$ are arbitrary paths connecting $A_{0}$ and $A_{1}$, Eq. (9) implies that the quantity $\int_{\gamma}\delta Q/T$ has a unique value independent of a path, which we denote by $\int_{0}^{1}\delta Q/T$. Therefore, we can define a state quantity $S$, which takes $S_{1}$ at the state $A_{1}$ as
@@
in a form independent of a path $\gamma$ by using the entropy. Equation (11) serves as the maximum work extracted from a system with the aid of a heat reservoir at fixed temperature (see Appendix A for a simple proof based on Ref. [8]). It should be noted again that the temperature $T$ in Eq. (10) is the system’s temperature, not the heat reservoir’s one as in Eq. (2). While Eq. (2) is usually obtained via the equality case of the Clausius inequality Eq. (3) that assumes thermal contact with multiple (or infinitely many) heat reservoirs with different temperatures, Eq. (10) has been obtained by directly applying the Kelvin-Planck principle assuming only a single heat reservoir at fixed temperature $T_{o}$. But, once Eq. (10) is established, we may interpret that it is equivalent to the Clausius entropy Eq. (2), as any path in Fig. 1 can be realized with a reversible process during which a system contacts multiple (or infinitely many) heat reservoirs with different temperatures, and the system’s temperature $T$ agrees with the temperature $T_{R}$ of each heat reservoir for such a reversible process.
@@
depending on the type of the $i$-th process. Here, $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is independent of $\gamma_{i}$ (i.e., determined only by the initial and final states), and note that $U(T^{(i+1)},V^{(i+1)})=U(T^{(i+1)},V^{(i)})$ in Eq. (13) because of $V^{(i+1)}=V^{(i)}$ (fixed volume) for the process (ii). These works (12) and (13) are obtained as follows: The work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is just given by the internal energy change $U^{(i)}-U^{(i+1)}=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})$. For the work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (13) for the process (ii), we operate a Carnot machine (heat engine, refrigerator, or heat pump) between the system at temperature $T$ and the heat reservoir at temperature $T_{o}$ until the system’s temperature $T$ initiating from $T^{(i)}$ reaches $T^{(i+1)}$ by keeping $V=V^{(i)}$. To raise the system’s temperature $T$, we may operate a Carnot heat engine for $T<T_{o}$ and a Carnot heat pump for $T_{o}\leq T$; to lower the system’s temperature $T$, we may operate a Carnot heat engine for $T_{o}<T$ and a Carnot refrigerator for $T\leq T_{o}$. For any case, the infinitesimal work $\delta W$ extracted by a Carnot machine operating between the system at $T$ and the heat reservoir at $T_{o}$ is commonly given by
@@
II.2 Derivation for simple systems
Figure 2: An arbitrary adiabatic process from one equilibrium state $A_{1}$ of a simple system to another equilibrium state $A_{2}$ (dashed curve), where $A_{1}$ and $A_{2}$ are created reversibley from a reference equilibrium state $A_{0}$ (solid curves). The Kelvin-Planck principle Eq. (5) is applied to the cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$.
@@ II.2 Derivation for simple systems
Let $A_{0}$ be a reference equilibrium state of the system specified with $(T_{0},V_{0})$. We create the equilibrium state $A_{1}$ from the reference equilibrium state $A_{0}$ reversibly by combining two thermodynamic processes with the aid of a single heat reservoir at fixed temperature $T_{o}$ considered in Sec. II.1 (Fig. 2). The work extracted from this process is given by Eq. (11):
@@
II.3 Derivation for compound systems
We show Eq. (1) from Eq. (5) for a compound system. Here, a compound system is composed of $M$ subsystems in equilibrium, where each subsystem is separated by other subsystems by insulated walls. The state of $j$-th subsystem is specified with $(T_{j},V_{j})$ ($j=1,\cdots,M$). The internal energy, entropy, and volume of the compound system are given by the sum of those of subsystems as $U=\sum_{j=1}^{M}U_{j}$, and $S=\sum_{j=1}^{M}S_{j}$, and $V=\sum_{j=1}^{M}V_{j}$, respectively.
@@ II.3 Derivation for compound systems
Subsequently, we create the state $A_{1}$ from a reference equilibrium state $A_{0}$ of a simple system specified with $(T_{0},V_{0})$. First, we change the volume from $V_{0}$ to $V_{1}$ by a reversible adiabatic process, resulting in the temperature $T_{0^{\prime}}$, where we denote the state $(T_{0^{\prime}},V_{1})$ by $A_{0^{\prime}}$. The reversible work extracted from this process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0^{\prime}}$ is given by
@@
III Discussion
III.1 Gyftopoulos-Beretta entropy
It should be noted that our formulation of entropy and its non-decrease in Sec. II may share some similarities with the approach by Gyftopoulos and Beretta (GB) [7] (see also Ref. [9]). By using notations in our formulation, their entropy is defined through the difference between works extracted from processes with and without the aid of a heat reservoir at fixed temperature $T_{o}$ [10]:
$S_{1}=S_{0}+\frac{W_{0\to 1}^{\rm rev}-W_{0\to 1}^{\rm ad/rev}}{T_{o}},\ S_{2}=S_{0}+\frac{W_{0\to 2}^{\rm rev}-W_{0\to 2}^{\rm ad/rev}}{T_{o}}.$ (27)
where $W_{0\to 1}^{\rm ad/rev}$ and $W_{0\to 2}^{\rm ad/rev}$ are the works extracted from the reversible adiabatic processes. This definition is equivalent to Eq. (10) and is derived from it with the use of Eq. (6) by noting $W_{0\to 1}^{\rm ad/rev}=U_{0}-U_{1}$ for a reversible adiabatic process, but the one focusing on work rather than heat. This is a natural definition as Eq. (11) suggests that the entropy serves to determine the reversible work as the maximum work extracted from a system with the aid of a heat reservoir at temperature $T_{o}$ [8]. Moreover, it should be noted that the GB entropy is referred to as the same entropy as the one derived by Lieb and Yngvason in their axiomatic thermodynamics [8] (see also Ref. [11] for a related discussion).

### v7

@@ Abstract
We derive the law of entropy non-decrease directly from the Kelvin-Planck principle for simple and compound systems without using the Clausius inequality. A key of the derivation is a new formulation of entropy in terms of work by a Carnot machine operating between a system and a single heat reservoir at fixed temperature, which is equivalent to Clausius entropy based on heat and Gyftopoulos-Beretta entropy based on work. We also show that we may characterize entropy as an extra thermodynamic cost that needs to be paid to create nonuniformity in the system.
1 Introduction
How information of irreversibility is encoded into entropy as a state quantity? This would be one of the most difficult points for those who learn thermodynamics first. This is also true for physicists today, and the famous paper by Lieb and Yngvason was aimed to elucidate the nature of entropy in terms of adiabatic accessibility in their axiomatic approach to thermodynamics [1, 2, 3], where neither statistical mechanics nor heat engines are used [3] (see also Ref. [4]).
Here, let us recall a conventional approach to entropy. Consider a thermally isolated system in an equilibrium state $A_{1}$. After an arbitrary adiabatic process that changes the system’s state to another equilibrium state $A_{2}$, the system’s entropy $S$ never decreases before and after the change:
@@ 1 Introduction
respectively, where $S_{0}$ is the entropy at a reference state $A_{0}$, $\delta Q$ denotes heat supplied to the system, and $T_{R}$ denotes the temperature of a heat reservoir. During the process from $A_{0}$ to $A_{1}$, the temperature $T_{R}$ of the heat reservoir that contacts the system may change. Equation (2) is the formula derived by Clausius. Usually, in standard textbooks on thermodynamics such as Ref. [5], after introducing Carnot cycle and Carnot efficiency as the maximum efficiency of heat engines, Eq. (1) is derived by using the Clausius inequality for a cyclic process:
@@
where $W_{\rm cyc}$ denotes the work extracted from a cycle starting from an equilibrium state, and the equality applies to reversible cycles. Deriving the Clausius inequality from the Kelvin-Planck principle relies on a complicated thermodynamic process that tactically involves multiple (or infinitely many) heat reservoirs at different temperatures and multiple (or infinitely many) Carnot heat engines [5]. This is because the Kelvin-Planck principle is originally applied to a process with a single heat reservoir at fixed temperature, and a complicated setup is necessary for it to be applied to a case with multiple (or infinitely many) heat reservoirs. Moreover, the law of entropy non-decrease Eq. (1) is derived as a special case of the Clausius inequality Eq. (3) (i.e., the case of $\delta Q=0$ in Eq. (4)) and it is thus derived indirectly from the Kelvin-Planck principle.
@@
In this paper, we show that we can derive the law of entropy non-decrease Eq. (1) directly from the Kelvin-Planck principle Eq. (5) without using the Clausius inequality Eq. (3). To this end, we give a new formulation of entropy of a substance in terms of work of a Carnot machine operating between the substance and a single heat reservoir at fixed temperature. Because we will assume only a single heat reservoir at fixed temperature in our derivation, it apparently makes it possible to apply the Kelvin-Planck principle to show the law of entropy non-decrease. We discuss the relationship between our entropy and the Clausius entropy (2) based on heat or the Gyftopoulos-Beretta entropy [6] based on work. Moreover, we also show that entropy may be characterized by an extra thermodynamic cost that needs to be paid to create nonuniformity in the system, which makes it possible to distinguish between two macroscopic systems of the same internal energy and volume.
2 Main results
@@
We derive the law of entropy non-decrease Eq. (1) from the Kelvin-Planck principle Eq. (5). After the formulation of our entropy in terms of work by a Carnot machine in Sec. 2.1, we give the derivation for simple systems in Sec 2.2 and compound systems in Sec. 2.3. Here, a simple system is a system in an equilibrium state that can be specified with the system’s temperature $T$ and volume $V$. A compound system is a system composed of simple systems.
2.1 Formulation of entropy as state quantity
@@
Figure 1: A reversible path $\gamma$ from $A_{0}$ to $A_{1}$ and a reversible path $\gamma^{\prime}$ from $A_{1}$ to $A_{0}$ on the temperature-volume (TTVV) plane. They consist of the reversible adiabatic processes (arrows along the adiabatic curves) and the reversible temperature-changing processes with a Carnot machine operating between the system with its volume being kept fixed and a heat reservoir at temperature $T_{o}$ (vertical arrows).
@@
We introduce our entropy as a state quantity in the following way. Consider a simple system initially in an equilibrium state $A_{0}$ specified with $(T_{0},V_{0})$ with $T_{0}$ and $V_{0}$ the system’s initial temperature and volume, respectively. Consider a reversible path $\gamma$ from the equilibrium state $A_{0}$ to another equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$ and a reversible path $\gamma^{\prime}$ from the state $A_{1}$ to $A_{0}$ (Fig. 1). Any reversible path $\gamma$ from the equilibrium state $A_{0}$ to another one $A_{1}$ can be realized by a combination of (i) a reversible adiabatic process (the arrow along the adiabatic curve in Fig. 1) and (ii) a reversible temperature-changing process with a Carnot machine operating between the system with its volume being kept constant and a single heat reservoir at fixed temperature $T_{o}$ to raise or lower the system’s temperature (the vertical arrow in Fig. 1) by regarding the system as a finite-sized heat reservoir [7]; any reversible path $\gamma$ may be decomposed into $N$-reversible processes $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$ with $\gamma_{i}$ being the $i$-th reversible process, which is either of the above two thermodynamic processes (i) and (ii). In particular, when the path is continuous, it may be realized by a combination of the processes (i) and (ii) by taking the limit of $N\to\infty$. It should be noted that there is no need for the system’s temperature to match the fixed temperature $T_{o}$ of the heat reservoir during the process including the initial and final states $A_{0}$ and $A_{1}$.
@@
where $\gamma^{\prime}$ denotes the reversed path of $\gamma^{\prime}$. Because $\gamma$ and $\gamma^{\prime}$ are arbitrary paths connecting $A_{0}$ and $A_{1}$, Eq. (9) implies that the quantity $\int_{\gamma}\delta Q/T$ has a unique value independent of a path, which we denote by $\int_{0}^{1}\delta Q/T$. Therefore, we can define a state quantity $S$, which takes $S_{1}$ at the state $A_{1}$ as
@@
in a form independent of a path $\gamma$ by using the entropy. Equation (11) serves as the maximum work extracted from a system with the aid of a heat reservoir at fixed temperature (see Appendix A for a simple proof based on Ref. [8]). It should be noted again that the temperature $T$ in Eq. (10) is the system’s temperature, not the heat reservoir’s one such as in Eq. (2). While Eq. (2) is usually obtained via the equality case of the Clausius inequality Eq. (3) that assumes thermal contact with multiple (or infinitely many) heat reservoirs with different temperatures, Eq. (10) has been obtained by directly applying the equality case of the Kelvin-Planck principle assuming only a single heat reservoir at fixed temperature $T_{o}$. But, once Eq. (10) is established, we may interpret that it is equivalent to the Clausius entropy Eq. (2), as any path in Fig. 1 can be realized with a reversible process during which a system contacts multiple (or infinitely many) heat reservoirs with different temperatures, and the system’s temperature $T$ agrees with the temperature $T_{R}$ of each heat reservoir for such a reversible process.
@@
depending on the type of the $i$-th process. Here, $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is independent of $\gamma_{i}$ (i.e., determined only by the initial and final system’s states), and note that $U(T^{(i+1)},V^{(i+1)})=U(T^{(i+1)},V^{(i)})$ in Eq. (13) because of $V^{(i+1)}=V^{(i)}$ (fixed volume) for the process (ii). These works (12) and (13) are obtained as follows: The work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is just given by the internal energy change $U^{(i)}-U^{(i+1)}=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})$. For the work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (13) for the process (ii), we operate a Carnot machine (heat engine, refrigerator, or heat pump) between the system at temperature $T$ and the heat reservoir at temperature $T_{o}$ until the system’s temperature $T$ initiating from $T^{(i)}$ reaches $T^{(i+1)}$ by keeping $V=V^{(i)}$. To raise the system’s temperature $T$, we may operate a Carnot heat engine for $T<T_{o}$ and a Carnot heat pump for $T_{o}\leq T$; to lower the system’s temperature $T$, we may operate a Carnot heat engine for $T_{o}<T$ and a Carnot refrigerator for $T\leq T_{o}$. For any case, the infinitesimal work $\delta W$ extracted by a Carnot machine operating between the system at $T$ and the heat reservoir at $T_{o}$ is commonly given by
@@
2.2 Derivation for simple systems
Figure 2: An arbitrary adiabatic process from one equilibrium state $A_{1}$ of a simple system to another equilibrium state $A_{2}$ (dashed curve), where $A_{1}$ and $A_{2}$ are created reversibly from a reference equilibrium state $A_{0}$ (solid curves). The Kelvin-Planck principle Eq. (5) is applied to the cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$.
@@ 2.2 Derivation for simple systems
Let $A_{0}$ be a reference equilibrium state of the system specified with $(T_{0},V_{0})$. We create the equilibrium state $A_{1}$ from the reference equilibrium state $A_{0}$ reversibly by combining two thermodynamic processes (i) and (ii) with the aid of a single heat reservoir at fixed temperature $T_{o}$ considered in Sec. 2.1 (Fig. 2). The work extracted from this process is given by Eq. (11):
@@
2.3 Derivation for compound systems
We show Eq. (1) from Eq. (5) for a compound system. Here, a compound system is composed of $M$ subsystems in equilibrium with different temperatures, where each subsystem is separated by other subsystems by insulated walls. So, a compound system may be regarded a nonuniform system. The state of $j$-th subsystem is specified with $(T_{j},V_{j})$ ($j=1,\cdots,M$). The internal energy $U$, entropy $S$, and volume $V$ of the compound system are given by the sum of those of subsystems as $U=\sum_{j=1}^{M}U_{j}$, and $S=\sum_{j=1}^{M}S_{j}$, and $V=\sum_{j=1}^{M}V_{j}$, respectively.
@@ 2.3 Derivation for compound systems
Subsequently, we create the state $A_{1}$ of the compound system from a reference equilibrium state $A_{0}$ of a simple system specified with $(T_{0},V_{0})$. First, we change the volume from $V_{0}$ to $V_{1}$ by a reversible adiabatic process, resulting in the temperature $T_{0^{\prime}}$, where we denote the state $(T_{0^{\prime}},V_{1})$ by $A_{0^{\prime}}$. The reversible work extracted from this process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0^{\prime}}$ is given by
@@
3 Discussion
3.1 Gyftopoulos-Beretta entropy
It should be noted that our formulation of entropy and its non-decrease in Sec. 2 may share some similarities with the approach by Gyftopoulos and Beretta (GB) [6] (see also Ref. [9]). By using notations in our formulation, their entropy is defined through the difference between works extracted from processes with and without the aid of a heat reservoir at fixed temperature $T_{o}$ [10]:
$S_{1}=S_{0}+\frac{W_{0\to 1}^{\rm rev}-(U_{0}-U_{1})}{T_{o}},\ S_{2}=S_{0}+\frac{W_{0\to 2}^{\rm rev}-(U_{0}-U_{2})}{T_{o}}.$ (27)
where we used $W_{0\to 1}^{\rm ad}=U_{0}-U_{1}$ and $W_{0\to 2}^{\rm ad}=U_{0}-U_{2}$ for the process without the aid of a heat reservoir. This definition is equivalent to Eq. (10) and is derived from it with the use of Eq. (6), but the one focusing on work rather than heat. This is a natural definition as Eq. (11) suggests that the entropy serves to determine the reversible work as the maximum work extracted from a system with the aid of a heat reservoir at temperature $T_{o}$ [8]. Moreover, it should be noted that the GB entropy is referred to as the same entropy as the one derived by Lieb and Yngvason in their axiomatic thermodynamics [8] (see also Ref. [11] for a related discussion).